# DCAD-2000: A Multilingual Dataset across 2000+ Languages with Data Cleaning as Anomaly Detection

### v2

@@ Abstract
The rapid development of multilingual large language models (LLMs) highlights the need for high-quality, diverse, and clean multilingual datasets. In this paper, we introduce DCAD-2000 (Data Cleaning as Anomaly Detection), a large-scale multilingual corpus built using newly extracted Common Crawl data and existing multilingual datasets. DCAD-2000 includes over 2,282 languages, 46.72TB of data, and 8.63 billion documents, spanning 155 high- and medium-resource languages and 159 writing scripts. To overcome the limitations of current data cleaning methods, which rely on manual heuristic thresholds, we propose reframing data cleaning as an anomaly detection task. This dynamic filtering approach significantly enhances data quality by identifying and removing noisy or anomalous content. We evaluate the quality of DCAD-2000 on the FineTask benchmark, demonstrating substantial improvements in multilingual dataset quality and task performance. [^1] We release DCAD-2000 at: https://huggingface.co/datasets/openbmb/DCAD-2000 and data cleaning pipeline at https://github.com/yl-shen/DCAD-2000.
DCAD-2000: A Multilingual Dataset across 2000+ Languages with Data Cleaning as Anomaly Detection
Yingli Shen1thanks: Equal contribution., Wen Lai2,3∗, Shuo Wang1, Xueren Zhang4, Kangyang Luo1 Alexander Fraser2,3, Maosong Sun1thanks: Corresponding author. 1 Tsinghua University, Beijing, China 2 Technical University of Munich, Germany 3 Munich Center for Machine Learning, Germany 4 Modelbest Inc., China
@@ 1 Introduction
Large language models (LLMs) have achieved great progress on a variety of NLP tasks by leveraging vast amounts of training data Minaee et al. (2024). However, their performance remains heavily biased towards high-resource languages (e.g., English), while low-resource languages are still left behind Zhu et al. (2024); Huang et al. (2024). To improve the multilingual capabilities of LLMs, a common strategy is to incorporate large amounts of non-English data, either by continue pretraining Lai et al. (2024) or by instruction tuning in multilingual settings Üstün et al. (2024). Therefore, constructing large-scale, high-quality multilingual datasets is crucial for enhancing the multilingual performance of LLMs.
@@
mC4 Raffel et al. (2020) CC-MAIN-2020-34 101 0 43 52 6
OSCAR 23.01 Abadji et al. (2022) CC-MAIN-2022-49 153 6 42 25 80
Glot500 Imani et al. (2023) CC-MAIN-2020-34 511 0 108 79 324
CulturaX Nguyen et al. (2024) CC-MAIN-2022-49 167 11 47 27 82
Madlad-400 Kudugunta et al. (2024) CC-MAIN-2022-33 419 7 46 39 327
MaLA Ji et al. (2024) CC-MAIN-2022-49 939 1 125 78 735
Glotcc Kargaran et al. (2024) CC-MAIN-2023-50 1331 0 10 52 1269
HPLT-v1.2 de Gibert et al. (2024) CC-MAIN-2022-40 191 12 53 38 88
Fineweb-2 Penedo et al. (2024b) CC-MAIN-2024-18 1915 10 62 49 1794
DCAD-2000 CC-MAIN-2024-46 2282 13 142 124 2003
Table 1: Comparison of multilingual datasets constructed from Common Crawl (CC) and our constructed DCAD-2000, focusing on the latest CC version used, the total number of languages supported, distribution across resource categories (high, medium, low, very low), and training readiness. The CC version marked with underline indicates an inferred version due to the lack of explicit specification in the original paper. The “Training-Ready” column indicates whether the dataset is ready for training LLMs without requiring further data cleaning.
@@
Recent efforts to construct multilingual datasets have built several notable resources, including CulturaX Nguyen et al. (2024), HPLT de Gibert et al. (2024), Madlad-400 Kudugunta et al. (2024), MaLA Lin et al. (2024), and Glotcc Kargaran et al. (2024), which cover 167, 191, 419, 939, and 1,331 languages, respectively. While these datasets have made significant contributions, they exhibit three major limitations, as summarized in Table 1: (1) Outdated data sources: These datasets primarily rely on older versions of Common Crawl [^2] https://commoncrawl.org, which results in outdated knowledge and an elevated risk of hallucination Huang et al. (2023). (2) Limited coverage of high- and medium-resource languages [^3] We follow the criteria from Goyal et al. (2022) to categorize languages: High: $>100M$; Medium: $(1M,100M)$; Low: $(100K,1M)$; Very Low: $<100K$.: For instance, Fineweb-2 Penedo et al. (2024b), which supports 1,915 languages, includes data from only 10 high-resource languages and 62 medium-resource languages. (3) Insufficient data cleaning: Despite being cleaned, recent studies Dou et al. (2024); Zhang et al. (2024) indicate that these datasets still contain a significant amount of noise, which makes them difficult to directly employ in training multilingual LLMs. For instance, Dou et al. (2024) found that 31.11% of the data in Madlad-400 could still be further removed by advanced data cleaning.
@@
Traditional data cleaning workflows Albalak et al. (2024) often rely on document-level features (e.g., language identification; Kargaran et al., 2023) and fixed thresholds to filter out low-quality data. However, this approach struggles with cross-lingual consistency due to feature distribution differences. For example [^4] Please refer to Appendix A for more details., the variation in average word counts across languages, datasets, and shards demonstrates the limitations of heuristic threshold-based data cleaning methods. Notably, while Fineweb-2 fine-tunes thresholds for more than 1,000 languages, this process is computationally intensive and time-consuming.
@@
To address these challenges, we introduce DCAD-2000, a new large-scale, high-quality multilingual dataset that can be directly applied to LLM training. DCAD-2000 covers 2282 languages (155 high/medium languages), incorporating the latest Common Crawl data (November 2024; CC-MAIN-2024-46) and existing multilingual datasets. Additionally, we propose a novel language-agnostic data cleaning approach that treats data cleaning as an anomaly detection Su et al. (2024) problem, distinguishing it from traditional threshold-based methods Laurençon et al. (2022); Penedo et al. (2024b). Our approach extracts eight statistical features from each document to evaluate quality, including metrics like language identification score, word repetition, special character ratio, and perplexity score. Anomaly detection algorithms dynamically identify and remove outliers by recognizing deviations from typical document quality metrics, ensuring consistent and language-agnostic filtering.
We provide a comprehensive analysis of DCAD-2000 (Section 4), highlighting its diverse document distribution, broad geographical and script coverage, as well as resource categorization across languages, with a particular focus on supporting both high-resource and underrepresented languages. The dataset’s extensive multilingual and cross-script representation is further validated through evaluation on the FineTask benchmark Penedo et al. (2024b), where LLMs trained on DCAD-2000 consistently outperform those trained on other multilingual datasets. Additionally, our anomaly detection-based data cleaning method demonstrates significant improvements in model performance, ensuring high-quality, noise-reduced data for large-scale multilingual model training.
@@
In summary, we make the following contributions: (1) We propose a novel data cleaning framework that frames the task as anomaly detection, offering a language-agnostic and adaptive solution without manual threshold tuning. (2) We release DCAD-2000, a comprehensive multilingual dataset covering over 2,282 languages, containing 8.63B of documents, 46.72TB of disk size and 159 writing scripts with metadata annotations, making it suitable for a wide range of downstream NLP tasks. (3) Extensive evaluation experiments on FineTask benchmark demonstrate that DCAD-2000 consistently outperforms existing multilingual corpora, achieving higher normalized accuracy across multiple languages and NLP tasks.
@@ 2 Related Work
2.1 Multilingual Dataset for Pretraining
Enhancing the multilingual capabilities of LLMs often involves continuing pretraining on large-scale multilingual datasets Lin et al. (2024); Ji et al. (2024). These datasets can be broadly categorized into curated corpora, domain-specific corpora, and web-crawled corpora.
Curated Corpora. Curated datasets are carefully gathered by experts from high-quality sources such as books Laurençon et al. (2022), academic publications Clement et al. (2019), and encyclopedia entries Brown (2011); Lehmann et al. (2015); Kuo et al. (2024). While providing highly accurate and reliable content, these datasets often cover only a small subset of languages, which limits their utility for training multilingual models.
Domain-Specific Corpora. In addition to general-domain data, fine-tuning LLMs on domain-specific datasets is crucial for improving performance in specialized domains like finance Zhang and Yang (2023), healthcare Wu et al. (2024), legal Colombo et al. (2024), and education Xu et al. (2024); Lozhkov et al. (2024). However, these datasets often narrow in scale and language diversity, typically supporting only a small group of high-resource languages, which limits their broader applicability.
Web-Crawled Corpora. Web-crawled datasets, particularly those derived from Common Crawl, provide large-scale multilingual coverage by leveraging an open repository of over 250 billion web pages. These datasets include mC4 Raffel et al. (2020), CC-100 Conneau et al. (2020), OSCAR Abadji et al. (2022), Glotcc Kargaran et al. (2024), Fineweb Penedo et al. (2024a), and Fineweb-2 Penedo et al. (2024b). However, these datasets typically only have basic data cleaning applied, and they still contain a significant amount of noisy or low-quality text, which makes them unsuitable for direct use in LLMs training. Moreover, the imbalance in language distribution is another issue, with particularly low coverage of high-resource and medium-resource languages.
2.2 Data Cleaning
Data cleaning is an essential step in preparing high-quality datasets for training robust LLMs. It involves filtering noisy, irrelevant, or harmful content and can be broadly classified into model-based and heuristic-based approaches Liu et al. (2024).
@@ 2.2 Data Cleaning
Model-Based Methods. Model-based approaches employ classifiers or LLMs to distinguish between high-quality and low-quality data. For instance, content safety models Li et al. (2024) filter out explicit or gambling-related content, while quality classifiers remove low-relevance text Jiang et al. (2024). LLM-based methods focus on generating prompts for cleaning Narayan et al. (2022) or integrating error detection and correction into the pipeline Chen et al. (2023); Ni et al. (2024). Although these methods are effective on small-scale datasets, they face scalability challenges due to the high computational cost of processing large-scale data. In particular, training a separate classifier for each language in a multilingual dataset is impractical.
Heuristic-Based Methods. Heuristic approaches apply predefined rules to filter content at both document and sentence levels. At the document level, strategies include filtering by language identification scores or scoring documents with language models Laurençon et al. (2022); Nguyen et al. (2024). At the sentence level, rules are applied to remove incomplete or irrelevant content, such as HTML tags or excessively short sentences Raffel et al. (2020); Abadji et al. (2022). Although computationally efficient, heuristic methods often require manual tuning and struggle to adapt to the diverse characteristics of multilingual corpora.
@@ 3 DCAD-2000
To overcome the limitations of existing multilingual datasets, we propose DCAD-2000, a large-scale, high-quality multilingual dataset constructed by integrating data from latest version of Common Crawl and existing multilingual datasets (Section 3.1). This dataset is cleaned using our proposed framework, which treats data cleaning as an anomaly detection problem (Section 3.2). The construction of DCAD-2000 is supported by robust computational resources, as detailed in Section 3.3.
@@ 3.1 Data Collection
To ensure comprehensiveness and robustness in multilingual data representation, DCAD-2000 integrates data from four main sources: MaLA, Fineweb, Fineweb-2, and newly extracted Common Crawl data. Each source is selected based on its unique contribution to multilingual coverage, data quality, and freshness, with careful consideration of dataset complementarity to avoid redundancy. Specifically, MaLA and Fineweb-2 are prioritized due to their broad language coverage and high-quality curation, which complements other widely used datasets like mC4 Raffel et al. (2020) and OSCAR Abadji et al. (2022).
MaLA Corpus Ji et al. (2024). The MaLA corpus covers 939 languages, aggregating data from diverse sources including Bloom Leong et al. (2022), CC100 Conneau et al. (2020), Glot500 Imani et al. (2023), among others. Deduplication is performed using MinHashLSH Broder et al. (1998), which is particularly effective in removing near-duplicate entries that often arise from common web sources. Language codes are based on ISO 639-3555https://en.wikipedia.org/wiki/ISO_639-3 standards, and language-specific scripts are supported by GlotScript [^6] https://github.com/cisnlp/GlotScript.
Fineweb Corpus Penedo et al. (2024a). Fineweb is a high-quality English web dataset extracted from Common Crawl, consisting of over 15 trillion tokens and updated monthly. Data cleaning and deduplication are performed using the Datatrove library. [^7] https://github.com/huggingface/datatrove For DCAD-2000, we incorporate data from the November 2024 release (CC-MAIN-2024-46) to ensure freshness and up-to-date relevance of the data.
Fineweb-2 Corpus Penedo et al. (2024b). Fineweb-2 expands Fineweb to include multilingual data, covering 1,915 languages. It processes 96 Common Crawl dumps from 2013 (CC-MAIN-2013-20) to April 2024 (CC-MAIN-2024-20). The deduplication process within Fineweb-2 is similarly handled using the Datatrove library, ensuring the exclusion of redundant entries and maintaining high-quality multilingual coverage.
@@ 3.2 Data Cleaning as Anomaly Detection
Traditional data cleaning methods rely on fixed thresholds for document-level features, making them less adaptable to the diversity of multilingual data. To address this, we propose a novel framework that formulates data cleaning as an anomaly detection task, which involves the feature extraction (Section 3.2.1) and anomaly detection (Section 3.2.2).
@@ 3.2.1 Feature Extraction
Inspired by Laurençon et al. (2022) and Nguyen et al. (2024), we extract eight statistical features from each document to evaluate text quality. Each feature is selected for its ability to capture important characteristics of the text, contributing to robust anomaly detection. Let $t$ represent a document; the extracted features are: (1) Number of Words, $n_{w}(t)$: Total token count after tokenization. (2) Character Repetition Ratio, $r_{c}(t)$: Fraction of repeated character sequences, highlighting noise or encoding errors. (3) Word Repetition Ratio, $r_{w}(t)$: Proportion of repeated words, indicating redundancy or unnatural text. (4) Special Characters Ratio, $r_{s}(t)$: Fraction of special characters based on language-specific lists from Laurençon et al. (2022). These lists include emojis, whitespace types, numbers, and punctuation. (5) Stopwords Ratio, $r_{\text{stop}}(t)$: Proportion of stopwords using Fineweb-2’s language-specific stopword lists. High or low ratios may indicate text quality issues. (6) Flagged Words Ratio, $r_{\text{flag}}(t)$: Fraction of toxic or profane words from Toxicity-200 Costa-jussà et al. (2022) or public repositories [^9] https://github.com/thisandagain/washyourmouthoutwithsoap. (7) Language Identification (LID) Score, $s_{\text{lid}}(t)$: Confidence score from GlotLID Kargaran et al. (2023), supporting over 2,000 languages. Low scores may indicate misclassification or noise. (8) Perplexity Score, $s_{\text{ppl}}(t)$: Using KenLM Heafield (2011), we train a language model for each language on multilingual Wikipedia data (November 2023). For unsupported languages, a default perplexity of 500 is used.
@@
$\mathbf{x}$ $=\left[n_{w}(t),\,r_{c}(t),\,r_{w}(t),\,r_{s}(t),\,\right.$ (1)
$\left.r_{\text{stop}}(t),\,r_{\text{flag}}(t),\,s_{\text{lid}}(t),\,s_{\text{ppl}}(t)\right]^{\top}\in\mathbb{R}^{8}.$

### v3

@@ Abstract
The rapid development of multilingual large language models (LLMs) highlights the need for high-quality, diverse, and clean multilingual datasets. In this paper, we introduce DCAD-2000 (Data Cleaning as Anomaly Detection), a large-scale multilingual corpus built using newly extracted Common Crawl data and existing multilingual datasets. DCAD-2000 includes over 2,282 languages, 46.72TB of data, and 8.63 billion documents, spanning 155 high- and medium-resource languages and 159 writing scripts. To overcome the limitations of current data cleaning methods, which rely on manual heuristic thresholds, we propose reframing data cleaning as an anomaly detection task. This dynamic filtering approach significantly enhances data quality by identifying and removing noisy or anomalous content. By fine-tuning LLMs on DCAD-2000, we demonstrate the effectiveness of the data quality improvements, the robustness of the data cleaning pipeline, and enhanced performance for low-resource languages across multiple multilingual benchmarks.
[Uncaptioned image] Dataset: https://huggingface.co/datasets/openbmb/DCAD-2000
[Uncaptioned image] Pipeline: https://github.com/yl-shen/DCAD-2000
@@ 1 Introduction
Large language models (LLMs) have achieved great progress on a variety of NLP tasks by leveraging vast amounts of training data [1]. However, their performance remains heavily biased towards high-resource languages [2, 3]. To improve the multilingual capabilities of LLMs, a common strategy is to incorporate large amounts of non-English data, either by continue pretraining [4] or by instruction tuning in multilingual settings [5]. Therefore, constructing large-scale, high-quality multilingual datasets is crucial for enhancing the multilingual performance of LLMs.
Table 1: Comparison of multilingual datasets constructed from Common Crawl ($C$) and our constructed DCAD-2000, focusing on the latest $C$ version used, the total number of languages supported, distribution across resource categories (high, medium, low, very low), and training readiness. The $C$ version marked with underline indicates an inferred version due to the lack of explicit specification in the original paper. The “Training-Ready” column indicates whether the dataset is ready for training LLMs without requiring further data cleaning.
@@
mC4 [6] $C$-MAIN-2020-34 101 0 43 52 6
OSCAR 23.01 [7] $C$-MAIN-2022-49 153 6 42 25 80
Glot500 [8] $C$-MAIN-2020-34 511 0 108 79 324
CulturaX [9] $C$-MAIN-2022-49 167 11 47 27 82
Madlad-400 [10] $C$-MAIN-2022-33 419 7 46 39 327
MaLA [11] $C$-MAIN-2022-49 939 1 125 78 735
Glotcc [12] $C$-MAIN-2023-50 1331 0 10 52 1269
HPLT-v1.2 [13] $C$-MAIN-2022-40 191 12 53 38 88
Fineweb-2 [14] $C$-MAIN-2024-18 1915 10 62 49 1794
DCAD-2000 $C$-MAIN-2024-46 2282 13 142 124 2003
@@
Recent efforts to construct multilingual datasets have built several notable resources, including CulturaX [9], HPLT [13], Madlad-400 [10], MaLA [15], and Glotcc [12], which cover 167, 191, 419, 939, and 1,331 languages, respectively. While these datasets have made significant contributions, they exhibit three major limitations, as summarized in Table 1: (1) Outdated data sources: These datasets primarily rely on older versions of Common Crawl [^1] https://commoncrawl.org, which results in outdated knowledge and an elevated risk of hallucination [16]. (2) Limited coverage of high- and medium-resource languages [^2] We follow the criteria from Flores-101 [17] to categorize languages: High: >100M>100M; Medium: $(1M,100M)$; Low: $(100K,1M)$; Very Low: <100K<100K.: For instance, Fineweb-2 [14], which supports 1,915 languages, includes data from only 10 high-resource languages and 62 medium-resource languages. (3) Insufficient data cleaning: Despite being cleaned, recent studies [18, 19] indicate that these datasets still contain a significant amount of noise, which makes them difficult to directly employ in training multilingual LLMs. For instance, Sailor [18] founds that 31.11% of the data in Madlad-400 could still be further removed by advanced data cleaning.
@@
Traditional data cleaning workflows [20] often rely on document-level features (e.g., language identification; 21) and fixed thresholds to filter out low-quality data. However, this approach struggles with cross-lingual consistency due to feature distribution differences. For example [^1] Please refer to Appendix A for more details., the variation in average word counts across languages, datasets, and shards demonstrates the limitations of heuristic threshold-based data cleaning methods. Notably, while Fineweb-2 fine-tunes thresholds for more than 1,000 languages, this process is computationally intensive and time-consuming.
@@
To address these challenges, we introduce DCAD-2000, a new large-scale, high-quality multilingual dataset that can be directly applied to LLM training. DCAD-2000 covers 2282 languages (155 high/medium languages), incorporating the latest Common Crawl data (November 2024; CC-MAIN-2024-46) and existing multilingual datasets. Additionally, we propose a novel language-agnostic data cleaning approach that treats data cleaning as an anomaly detection [22] problem, distinguishing it from traditional threshold-based methods [14, 23]. Our approach extracts eight statistical features from each document to evaluate quality, including number of words, character/word repetition ratio, special character/word ratio, stopword ratio, flagged words ratio, language identification score and perplexity score. Anomaly detection algorithms dynamically identify and remove outliers by recognizing deviations from typical document quality metrics.
We conduct a comprehensive analysis of DCAD-2000 with respect to document distribution, linguistic and geographical characteristics, writing scripts, and resource classification (Section 4). By fine-tuning LLMs on DCAD-2000, we validate the effectiveness of its data quality and data cleaning pipeline. Furthermore, we demonstrate the superiority of DCAD-2000 across various language categories (high, medium, low and very low) in multiple multilingual benchmarks, including SIB-200 [24], Glot500 [8] and FLORES-200 [25] (Section 5).
@@
In summary, we make the following contributions: (1) We propose a novel data cleaning framework that frames the task as anomaly detection, offering a language-agnostic and adaptive solution without manual threshold tuning. (2) We release DCAD-2000, a comprehensive multilingual dataset covering over 2,282 languages, containing 8.63B of documents, 46.72TB of disk size and 159 writing scripts with metadata annotations. (3) Extensive evaluation across multiple multilingual benchmarks demonstrates the effectiveness of both the data quality and the data cleaning pipeline.
@@ 2 Related Work
Multilingual Dataset for Pretraining.
Enhancing the multilingual capabilities of LLMs often involves continuing pretraining on large-scale multilingual datasets [11, 15]. These datasets can be broadly categorized into curated corpora, domain-specific corpora, and web-crawled corpora. (I) Curated Corpora. Curated datasets are carefully gathered by experts from high-quality sources such as books [23], academic publications [26], and encyclopedia entries [27, 28, 29]. (II) Domain-Specific Corpora. In addition to general-domain data, fine-tuning LLMs on domain-specific datasets is crucial for improving performance in specialized domains like finance [30], healthcare [31], legal [32], and education [33, 34]. (III) Web-Crawled Corpora. Web-crawled datasets, particularly those derived from Common Crawl, provide large-scale multilingual coverage by leveraging an open repository of over 250 billion web pages. These datasets include mC4 [6], CC-100 [35], OSCAR [7], Glotcc [12], Fineweb [36], and Fineweb-2 [14]. While curated and domain-specific corpora offer high-quality content with limited language coverage, web-crawled corpora provide broader multilingual coverage but often suffer from noise and lower data quality [18, 19].
Data Cleaning.
@@
Data cleaning is an essential step in preparing high-quality datasets for training robust LLMs. It involves filtering noisy, irrelevant, or harmful content and can be broadly classified into model-based and heuristic-based approaches [37]. (I) Model-Based Methods. Model-based approaches employ classifiers or LLMs to distinguish between high-quality and low-quality data. For instance, content safety models [38] filter out explicit or gambling-related content, while quality classifiers remove low-relevance text [39]. LLM-based methods focus on generating prompts for cleaning [40] or integrating error detection and correction into the pipeline [41, 42]. (II) Heuristic-Based Methods. Heuristic approaches apply predefined rules to filter content at both document and sentence levels. At the document level, strategies include filtering by language identification scores or scoring documents with language models [9, 23]. At the sentence level, rules are applied to remove incomplete or irrelevant content, such as HTML tags or excessively short sentences [6, 7]. While model-based methods offer high precision but face scalability challenges, heuristic-based methods are more efficient yet less adaptable to diverse multilingual data.
@@ 3 DCAD-2000
To overcome the limitations of existing multilingual datasets, we introduce DCAD-2000, a large-scale, high-quality multilingual dataset constructed by integrating data from latest version of Common Crawl and existing multilingual datasets (Section 3.1). This dataset is cleaned using our proposed novel framework, which treats data cleaning as an anomaly detection problem (Section 3.2). The construction of DCAD-2000 is supported by robust computational resources, as detailed in Section 3.3.
@@ 3.1 Data Collection
To ensure comprehensive and robust multilingual data representation, DCAD-2000 integrates data from four main sources: MaLA, Fineweb, Fineweb-2, and newly extracted Common Crawl data. Each source is selected based on its unique contribution to multilingual coverage, data quality, and freshness, with careful consideration to complementarity to minimize redundancy. Specifically, MaLA and Fineweb-2 are prioritized due to their broad language coverage and high-quality curation, which complements other widely used datasets like mC4 [6] and OSCAR [7].
MaLA Corpus [11]. The MaLA corpus covers 939 languages, aggregating data from diverse sources including Bloom [43], CC100 [35], Glot500 [8], among others. Deduplication is performed using MinHashLSH [44], which is particularly effective in removing near-duplicate entries that often arise from common web sources. Language codes are based on ISO 639-3222https://en.wikipedia.org/wiki/ISO_639-3 standards, and language-specific scripts are supported by GlotScript [^3] https://github.com/cisnlp/GlotScript.
Fineweb Corpus [36]. Fineweb is a high-quality English web dataset extracted from Common Crawl, consisting of over 15 trillion tokens and updated monthly. Data cleaning and deduplication are performed using the Datatrove library. [^4] https://github.com/huggingface/datatrove For DCAD-2000, we incorporate data from the November 2024 release (CC-MAIN-2024-46) to ensure freshness and up-to-date relevance of the data.
Fineweb-2 Corpus [14]. Fineweb-2 expands Fineweb to include multilingual data, covering 1,915 languages. It processes 96 Common Crawl dumps from 2013 (CC-MAIN-2013-20) to April 2024 (CC-MAIN-2024-20). The deduplication process within Fineweb-2 is similarly handled using the Datatrove library, ensuring the exclusion of redundant entries and maintaining high-quality multilingual coverage.
@@ 3.2 Data Cleaning as Anomaly Detection
Traditional data cleaning methods rely on fixed thresholds for document-level features, making them less adaptable to the diversity of multilingual data. To address this, we propose a novel framework that formulates data cleaning as an anomaly detection task, which involves feature extraction (Section 3.2.1) and anomaly detection (Section 3.2.2).
@@ 3.2.1 Feature Extraction
Inspired by Roots [23] and CulturaX [9], we extract eight statistical features from each document to evaluate text quality. Each feature is selected for its ability to capture important characteristics of the text, contributing to robust anomaly detection. Let $t$ represent a document; the extracted features are: (1) Number of Words, $n_{w}(t)$: Total token count after tokenization. (2) Character Repetition Ratio, $r_{c}(t)$: Fraction of repeated character sequences, highlighting noise or encoding errors. (3) Word Repetition Ratio, $r_{w}(t)$: Proportion of repeated words, indicating redundancy or unnatural text. (4) Special Characters Ratio, $r_{s}(t)$: Fraction of special characters based on language-specific lists from ROOTs Corpus [23]. These lists include emojis, whitespace types, numbers, and punctuation. (5) Stopwords Ratio, $r_{\text{stop}}(t)$: Proportion of stopwords using Fineweb-2’s language-specific stopword lists. (6) Flagged Words Ratio, $r_{\text{flag}}(t)$: Fraction of toxic or profane words from Toxicity-200 [25] or public repositories [^6] https://github.com/thisandagain/washyourmouthoutwithsoap. (7) Language Identification (LID) Score, $s_{\text{lid}}(t)$: Confidence score from GlotLID [21], supporting over 2,000 languages. (8) Perplexity Score, $s_{\text{ppl}}(t)$: Using KenLM [45], we train a language model for each language on multilingual Wikipedia data (November 2023) [^7] For languages not covered by the trained models, a default perplexity score of 500 is assigned..
@@
$\mathbf{x}=\left[n_{w}(t),\,r_{c}(t),\,r_{w}(t),\,r_{s}(t),\,r_{\text{stop}}(t),\,r_{\text{flag}}(t),\,s_{\text{lid}}(t),\,s_{\text{ppl}}(t)\right]^{\top}\in\mathbb{R}^{8}.$ (1)